satellite73:

the initial temp of an beer is 70 degrees, but it in a 10 degree freezer 1/2 a minute later it is only 50 degrees

satellite73:

work with the difference in temperatures, that is what decays

satellite73:

70-10=?

Vrefela:

60

satellite73:

so initial difference is 60, like having (0,60) on the graph

Vrefela:

right i understand this part

satellite73:

then in half a minute, what is the difference in temps?

Vrefela:

10?

satellite73:

jeez read the damned problem

satellite73:

in the book i mean, just read what it says in the book

Vrefela:

50

satellite73:

NO KEEP READING

Vrefela:

i am

Vrefela:

15?

satellite73:

A thermometer is removed from a room where the temperature is 70° F and is taken outside, where the air temperature is 10° F. After one-half minute the ther- mometer reads 50° F.

satellite73:

the initial difference it 70 - 10 = 60 after half a minute the difference is 50 - 10 = ?

Vrefela:

OH

satellite73:

oh...

Vrefela:

I don't like the way they worded it

satellite73:

too bad we are not done

Vrefela:

ok

satellite73:

we have (0,60) and (.5,40) we need the exponential model for that

satellite73:

\[T=60e^{kt}\] where if \[t=.5, T=40\] solve \[40=60e^{.5k}\] for \[k\]

Vrefela:

ln(.2)/.5

Vrefela:

wait

Vrefela:

.66

satellite73:

\[40\div 60\neq .2\]

Vrefela:

ln(.66)/.5

satellite73:

ok good enough now be careful

satellite73:

those are the temperature DIFFERENCES the actual temp you have to add the 10 degree ambient temperature so it really should be \[E=60e^{-.118t}+10\]

Vrefela:

so because there is a difference between the actual and whatever they give you, you have to add the difference?

satellite73:

you have to add the ambient temp, it is the differences that decays

Vrefela:

oh ok

satellite73:

the temp of the heated object decays to the ambient temp

satellite73:

now if you want the diffeq solution, which is identical, start with \[\frac{dT}{dt}=k(T-10)\]

Vrefela:

ok

satellite73:

separate get \[\frac{dT}{T-10}=kdt\] so \[\int \frac{dT}{T-10}=\int kdt\] making \[\ln(T-10)=kt\] and therefore \[T-10=Ce^{kt}\]

satellite73:

but it is identical only with the +10 out at the end the initial difference it 60 so C = 60 and you still find k the same way

Vrefela:

T = Ce^(kt) + 10

Vrefela:

This is way easier than the other way

satellite73:

identical

Vrefela:

easier

satellite73:

identical

satellite73:

and now i am done try one yourself, see that they are not that hard i think the mixture ones might use an integrating factor, not sure they separate

Vrefela:

I will tomorrow. I have to get to bed

Vrefela:

Thank you so much!

satellite73:

read example in 1.3 of setting up the mixture then read example of that same problem in 3.1

satellite73:

good night!

satellite73:

yw

Vrefela:

okie dokie sounds good! Goodnight!